OpenStudy (anonymous):

determine the average rate of change of the following functions for the given intervals. f(x) = 2/x+2 from x to x+h

OpenStudy (anonymous):

\[f(x+h)-f(x)/(x+h-x)=f(x+h)-f(x)/h\]

OpenStudy (anonymous):

\[f(x+h)=2/(x+h+2)\]

OpenStudy (anonymous):

and what about the rest

OpenStudy (anonymous):

[2/(x+h) + 2 - 2/x - 2]/h =-2/x(x+h) that is the rate of change.